OpenStudy (anonymous):

is n=4a solution of the equation 16=3n+4?

OpenStudy (agreene):

Given: f(n)≡ 16 = 3n +4 Lemma 1: Determine, if n=4 is a solution of f(n). 16 = 3(4) + 4 16 = 12 + 4 16 = 16 [1] ∵ [1] ∨ f(4) | ∃ n ∈ f(n), ∴ Lemma 1 is true.